OpenStudy (anonymous):

given that \[B_{n} = \sum_{i=0}^{n}(2/n)(4 - ((4i^2) / (n^2)))\] is the Riemann sum of an integrable function f(r) on an interval I. Find f(r) and I then evaluate \[\lim_{n \rightarrow \infty}B_{n}\] by integrating f(r) on I.

OpenStudy (anonymous):

another cool question :)

OpenStudy (anonymous):

this one? \[\sum _{i=0}^n \left(\frac{2}{n}\left(\frac{4-4i^2}{n^2}\right)\right)\]

OpenStudy (anonymous):

nono the 4 should be another rational k?

OpenStudy (anonymous):

its 4 - (b / a) (prototype)

OpenStudy (anonymous):

\[B_n = \sum_{1=0}^{n} \frac{2}{n}* (4 - \frac{4*i^2}{n^2})\]Right?

OpenStudy (anonymous):

yes its right :D

OpenStudy (anonymous):

Now, what am I supposed to do here

OpenStudy (anonymous):

u should find f(r) and evaluate \[\lim_{n \rightarrow \infty}B _{n}\]

OpenStudy (anonymous):

Should I integrate the whole function \[B_n = \sum_{i=0}^{n} \frac{2}{n}* (4 - \frac{4*i^2}{n^2})\] from 0 to n ...to find \(f(r)\)

OpenStudy (anonymous):

hmm i don't know how can i solve this thats why i asked :D

OpenStudy (anonymous):

Maybe we need integrating it, as for evaluating function \(n \rightarrow \infty\) we need to have some value for \(i\) too I mean we need some thing to get in place of \(i\)

OpenStudy (anonymous):

ah yea we need

OpenStudy (anonymous):

or not o.O

OpenStudy (anonymous):

Ah I have a method ... \[B_n = \sum_{i=0}^{n} \frac{8}{n} - \frac{8*i^2 }{n^3}\] Now sum square of n natural numbers is \(\frac{n(n+1)(2n+1)}{6}\) So I think the \(B_n\) thing should become \[B_n = \frac{8}{n}*n - \frac{8}{n}*\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ah yea

OpenStudy (anonymous):

In this riemann sum, the width of each rectangle is \[\frac{2}{n}\] if 2 is the interval being evaluated. This would make the height of each rectangle equal\[f(i) = \left(2-\frac{2i}{n}\right)\left(2+\frac{2i}{n}\right)\] The actual x value of for each i value is related as \[x = i \times \frac{2}{n}\] so we can substitute that in the previous equation to find the original function being integrated: \[f(x) = (2-x)(2+x)\] You can then integrate it from 0 to 2\[\int\limits_{0}^{2}(4-x^{2})dx = 4x-\frac{x^{3}}{3} +C = 5.333... - 0 = 5.333... \] I guess this is right, but I'm not sure...

OpenStudy (anonymous):

\[B_n = \frac{8}{n}*n - \frac{8}{n^3}*\frac{n(n+1)(2n+1)}{6}\]

OpenStudy (anonymous):

I didn't included powers typo

OpenStudy (anonymous):

\[B_n = 8 - 8*\frac{2n^2 + 3n +1 }{6n^2 }\]

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} 8 - \frac{8}{6}*(2 + \frac{3}{n} +\frac{1}{n^2})\]

OpenStudy (anonymous):

\[\frac{1}{n} \rightarrow 0\]

OpenStudy (anonymous):

\[8 -\frac{8*2}{6}\]

OpenStudy (anonymous):

yeah I get the same thing as alexray

OpenStudy (anonymous):

5.33333333333333333333333333333333.........................

OpenStudy (anonymous):

thanks anyway :D i have like 200-300 questions like that lol

OpenStudy (anonymous):

More 3s!!

OpenStudy (anonymous):

lol @ alexray19

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Please do post more I want to solve them too

OpenStudy (anonymous):

kk give me some time, i post one more .D :D

OpenStudy (anonymous):

:D